Problem 1

A researcher sends a questionnaire to 6,000 individuals. She gets back 3,000 responses. The percentage of women among the respondents is 40%. Because of nonresponse, the fraction of women among recipients of the survey could be anything between (low) a)_____? and (high) b) ______?
Answer #1

Answer)

Total recipients = 6000

Response = 3000

40% of respondents are females = 49% of 3000 = 1200

Now we can gaurantee that 1200 are women

So, lower limit = 1200/6000 = 0.2

Now, if 40% are women then 60% are male

60% of 3000 = 1800

So, highest number of women can be 6000-1800 = 4200

Upper limit = 4200/6000 = 0.7

So,

A) 0.2

B) 0.7